A television sports commentator wants to estimate the proportion of citizens who​ "follow professional​ football." Complete parts​ (a) through​ (c).

​(a) What sample size should be obtained if he wants to be within 2 percentage points with 96​% confidence if he uses an estimate of 54​% obtained from a​ poll?The sample size is ___.​(Round up to the nearest​ integer.)

​(b) What sample size should be obtained if he wants to be within 2 percentage points with 96​% confidence if he does not use any prior​ estimates?The sample size is ___.​(Round up to the nearest​ integer.)

​(c) Why are the results from parts​ (a) and​ (b) so​ close?

A.The results are close because the margin of error 2​% is less than​ 5%.

B.The results are close because 0.54 (1−0.54)=0.2484 is very close to 0.25.

C.The results are close because the confidence 96​% s close to​ 100%.
Answer #1

Ans:

Margin of error=0.02

confidence level=0.96

alpha=1-0.96=0.04

alpha/2=0.02

critical z value=normsinv(0.02) or normsinv(0.98)=2.054

a)prior estimate p=0.54

Sample size required,n=2.054^2*0.54*(1-0.54)/0.02^2=2620

b)no prior estimate,so p=0.5

Sample size required,n=2.054^2*0.5*(1-0.5)/0.02^2=2637

c)

Option B is correct.

The results are close because 0.54 (1−0.54)=0.2484 is very close to 0.25.